# VinePPO: Unlocking RL Potential For LLM Reasoning Through Refined Credit Assignment

### v1

@@ Abstract
Large language models (LLMs) are increasingly applied to complex reasoning tasks that require executing several complex steps before receiving any reward. Properly assigning credit to these steps is essential for enhancing model performance. Proximal Policy Optimization (PPO), a state-of-the-art reinforcement learning (RL) algorithm used for LLM finetuning, employs value networks to tackle credit assignment. However, value networks face challenges in predicting the expected cumulative rewards accurately in complex reasoning tasks, often leading to high-variance updates and suboptimal performance. In this work, we systematically evaluate the efficacy of value networks and reveal their significant shortcomings in reasoning-heavy LLM tasks, showing that they barely outperform a random baseline when comparing alternative steps. To address this, we propose VinePPO, a straightforward approach that leverages the flexibility of language environments to compute unbiased Monte Carlo-based estimates, bypassing the need for large value networks. Our method consistently outperforms PPO and other RL-free baselines across MATH and GSM8K datasets with fewer gradient updates (up to 9x), less wall-clock time (up to 3.0x). These results emphasize the importance of accurate credit assignment in RL finetuning of LLM and demonstrate VinePPO’s potential as a superior alternative [^1] Code available at https://github.com/McGill-NLP/VinePPO.
@@ 1 Introduction
Large language models (LLMs) are increasingly used for tasks requiring complex reasoning, such as solving mathematical problems (OpenAI, 2024), navigating the web (Zhou et al., 2024), or editing large codebases (Jimenez et al., 2024). In these settings, LLMs often engage in extended reasoning steps, executing multiple actions to arrive at a solution. However, not all steps are equally important—some contribute significantly, while others are irrelevant or detrimental. For example, in Figure 1.a, only step $s_{2}$ provides a key insight. Indeed, most reasoning steps generated by a model do not affect the chance of it solving the problem (Figure 1.b). Identifying the contribution of each action is crucial for improving model performance. However, this is inherently difficult due to the significant delay between actions and their eventual effect. This issue, known as the credit assignment problem, is a core challenge in reinforcement learning (RL, Sutton and Barto 1998).
@@
Proximal Policy Optimization (PPO, Schulman et al. 2017; Ouyang et al. 2022), a state-of-the-art algorithm for RL-based finetuning of LLMs (Xu et al., 2024; Ivison et al., 2024), tackles credit assignment using a value network (or critic). This network, typically a separate model initialized from a pretrained checkpoint, is trained during PPO finetuning to estimate the expected cumulative rewards (or value) of an intermediate action. In Figure 1.b, an ideal value network would assign high value to step $s_{2}$ and subsequent steps, where the model predicted a critical action. PPO uses these value estimates to measure the advantage of each action and update the model accordingly.
Accurately modeling value—predicting future rewards from an incomplete response—requires the value network to understand both the space of correct solutions (the very task the policy model is trying to learn) and predict the model’s future behavior, both of which are inherently challenging. In fact, there are hints in the literature that standard PPO implementations for LLMs have inaccurate value estimations. Ahmadian et al. (2024) and Trung et al. (2024) find that value networks often serve best as just a baseline in policy gradient222setting the Generalized Advantage Estimation (GAE, Schulman et al. 2016) parameter $\lambda=1$. Shao et al. (2024) show that the value network can be replaced by averaging rewards of responses to a given problem without degradation in performance.
Prompt ($s_{0}$) $\hat{p}(\text{correct}|s_{:t})$
@@
$0.4$ Let $a$ and $b$ be nonzero real numbers such that $(2-7i)(a+bi)$ is pure imaginary. Find $\frac{a}{b}.$
Response
$s_{1}$$s_{1}$$0.4$ We can expand the left-hand side to get
$s_{2}$$s_{2}$$\mathbf{1.0}$ $(2-7i)(a+bi)=(2a+7b)+(-7a+2b)i.$
$s_{3}$$s_{3}$$1.0$ This is pure imaginary if and only if the real part is 00, i.e.
$s_{4}$$s_{4}$$1.0$ $02a+7b=$ 0 .
$s_{5}$$s_{5}$$1.0$ Then $a=-\frac{7}{2}b,$ so $\frac{a}{b}=-\frac{7}{2}.$
A Sample Response
$\hat{p}(\text{correct}|s_{:t+1})-\hat{p}(\text{correct}|s_{:t})$
@@ A Sample Response
Figure 1: (Left) A response generated by the model. The notation $\hat{p}(\text{correct}|s_{:t})$ represents the estimated probability of successfully solving the problem at step $t$. Here, only step $s_{2}$ is critical; after this, the model completes the solution correctly. (Right) The delta in probability of successful completion between response steps. Most steps show little or no advantage over the preceding step.
Since errors in value estimation can lead to poor credit assignment and negatively impact convergence and performance (Greensmith et al., 2001), a natural question to ask is: how accurately do value networks actually perform during LLM finetuning? If we could improve credit assignment, to what extent would it enhance LLM performance? While recent studies (Hwang et al., 2024; Setlur et al., 2024) have begun to highlight the importance of identifying incorrect reasoning steps and incorporating them via ad-hoc mechanisms in “RL-free” methods (Rafailov et al., 2023), the broader question of how improving credit assignment might boost RL fine-tuning for LLMs remains open.
In this work, we evaluate the standard PPO pipeline in mathematical reasoning tasks across various model sizes. We find that value networks consistently provide inaccurate estimates and struggle to rank alternative steps correctly, suggesting that current PPO finetuning approaches for LLMs operate without effective credit assignment. To address this issue and illustrate the effect of accurate credit assignment, we propose VinePPO (Figure 2). Instead of relying on value networks, VinePPO computes unbiased value estimates of intermediate states by using independent Monte Carlo (MC) samples and averaging their respective return. This straightforward modification to PPO takes advantage of a special property of the language environment: the ability to easily reset to any intermediate state along the trajectory.
VinePPO consistently outperforms standard PPO and “RL-free” baselines, especially on more challenging datasets. Despite its slower per-iteration speed, it reaches and surpasses PPO’s peak performance with fewer gradient updates, resulting in less wall-clock time and lower KL divergence from the base model. Our findings highlight the importance of precise credit assignment in LLM finetuning and establishes VinePPO as a straightforward alternative to value network-based approaches.
Our contributions are as follows:
  • We demonstrate the suboptimal credit assignment in standard PPO finetuning by analyzing the value network, showing that it provides inaccurate estimates of intermediate state values and barely outperforms a random baseline when ranking alternative steps (see Section 7 for details).
  • We propose VinePPO, introduced in Section 4, which takes advantage of the flexibility of language as an RL environment to compute unbiased value estimates, eliminating the need for large value networks and reducing memory requirements (up to 112GB for a 7B LLM).
@@
  • VinePPO highlights the significance of credit assignment: It outperforms PPO and other baselines, especially on more challenging datasets. It achieves PPO’s peak performance with fewer iterations (up to 9x), less wall-clock time (up to 3.0x), and better KL-divergence trade-off. See Section 6.
[…]
Figure 2: (a) PPO finetunes the model by adjusting action probabilities based on their advantage, which is primarily guided by the value network’s value estimates. (b) VinePPO modifies standard PPO and obtains values estimates by simply resetting to intermediate states and using MC samples.
@@ 2 Related Work
PPO, as applied in RL from Human Feedback (RLHF, Ouyang et al. 2022), pioneered RL finetuning of LLMs. However, its computational overhead and hyperparameter sensitivity led to the development of simpler alternatives. RL-free methods such as DPO (Rafailov et al., 2023) operate in a bandit setting, treating the entire response as a single action. Similarly, rejection sampling methods like RestEM (Singh et al., 2024) finetune on full high-reward responses. RLOO (Ahmadian et al., 2024) and GRPO (Shao et al., 2024) abandon PPO’s value network, instead using average reward from multiple samples as a baseline. Recent work has emphasized finer credit assignment, with Hwang et al. (2024) and Setlur et al. (2024) introducing MC-based methods to detect key errors in reasoning chains for use as ad-hoc mechanisms in DPO. Our work, by contrast, fully embraces the RL training, with the target of unlocking PPO’s potential. Parallel efforts have also focused on building better verifiers and reward models for per-step feedback, with recent attempts to automate their data collection using MC rollouts (Ma et al., 2023; Uesato et al., 2022; Luo et al., 2024; Wang et al., 2024). Our method is orthogonal to these methods, operating within PPO-based training to optimize a given reward, instead of designing new ones.
@@
Deep RL algorithms are typically categorized into value-based and policy-based methods. Policy-based methods like PPO usually employ critic networks for value prediction. An exception is the “Vine” variant of TRPO (Schulman et al., 2015), which uses MC samples for state value estimation. The authors, however, note that the Vine variant is limited to environments that allow intermediate state resets, rare in typical RL settings [^3] This is reflected in the design of Gym (Towers et al., 2024), which only allows resets to the initial state.. However, language generation – when formulated as RL environment – enables such intermediate reset capabilities. In domains with similar reset capabilities, such as Go and Chess, MC-heavy methods like AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) have emerged. AlphaGo’s architecture includes a policy, trained using expert moves and self-play, and a value network that predicts game outcomes. At inference, it employs tree search guided by MC rollouts and value network to select optimal moves. AlphaZero advances this approach by distilling MCTS outcomes into the policy. Recent works have adapted AlphaZero’s principles to LLMs, employing similar search techniques during inference to improve responses and during training to find better trajectories for distillation (Xie et al., 2024; Chen et al., 2024; Wan et al., 2024; Zhang et al., 2024; Hao et al., 2023). While this is a promising direction, our method is not an MCTS approach; it uses MC samples solely for value estimation during PPO training to improve credit assignment.
Figure 3: VinePPO outperforms standard PPO and other RL-free baselines on Pass@1 performance on MATH and GSM8K datasets, while also exhibiting scalability across different model sizes.
@@ 3 Background
We focus on the RL tuning phase in the RLHF pipeline, following Ouyang et al. (2022); Shao et al. (2024). In this section, we provide an overview of actor-critic finetuning as implemented in PPO.
@@
where $\mathcal{D}$ is the dataset of prompts, $\mathcal{R}({\bm{x}};{\bm{y}})$ is the complete sequence-level reward function, and $\beta$ controls the strength of the KL penalty. Note that the policy $\pi_{\theta}$ is initialized from $\pi_{\mathrm{ref}}$.
@@
${\bm{g}}_{\mathrm{pg}}=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})A(s_{t},a_{t})\right],\quad\text{where}\ s_{t}={\bm{x}};{\bm{y}}_{<t},\ a_{t}=y_{t},$ (2)
where $A(s_{t},a_{t})$ is the advantage function. If $A(s_{t},a_{t})>0$, ${\bm{g}}_{\mathrm{pg}}$ will increase the probability of action $a_{t}$ in state $s_{t}$, and decrease it when $A(s_{t},a_{t})<0$. Intuitively, the advantage function quantifies how much better action $a_{t}$ is compared to average actions taken in state $s_{t}$ under the policy. Formally, it is defined as:
$A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})=r_{t}+\gamma V(s_{t+1})-V(s_{t}),$ (3)
@@
In practice, the advantage $A(s_{t},a_{t})$ is not known beforehand and is typically estimated by first using a value network $\hat{V}_{\phi}$ to approximate the true value function $V(s_{t})$, then substituting the learned values into Equation 3 or alternative methods like GAE (Schulman et al., 2016). The value network is parameterized by $\phi$ and trained alongside the policy network $\pi_{\theta}$. The training objective for the value network minimizes the mean squared error between the predicted value and the empirical return:
@@
where $G_{t}=\sum_{t^{\prime}=t}^{T-1}r_{t^{\prime}}$ is the empirical return from state $s_{t}$. PPO uses the same objective for $\hat{V}_{\phi}$ but enhances stability by applying clipping during training (see Section A.1 for details). In RL-tuning of LLMs, the value network is often initialized using the initial SFT policy $\pi_{\mathrm{ref}}$ (or the reward model when available), with the language modeling head swapped out for a scalar head to predict values (Zheng et al., 2023). This setup leverages the prior knowledge of the pretrained model.
Figure 4: Impact of number of sampled trajectories $K$ for estimating $\hat{V}_{\mathrm{MC}}(s_{t})$, evaluated on RhoMath 1.1B models. Increasing the number of rollouts improves task performance consistently.

### v2

# VinePPO: Refining Credit Assignment in RL Training of LLMs

@@ Abstract
Large language models (LLMs) are increasingly applied to complex reasoning tasks that require executing several complex steps before receiving any reward. Properly assigning credit to these steps is essential for enhancing model performance. Proximal Policy Optimization (PPO), a common reinforcement learning (RL) algorithm used for LLM finetuning, employs value networks to tackle credit assignment. However, recent approaches achieve strong results without it, raising questions about the efficacy of value networks in practice. In this work, we systematically evaluate the efficacy of value networks and reveal their significant shortcomings in reasoning-heavy LLM tasks, showing that they often produce poor estimate of expected return and barely outperform a random baseline when comparing alternative steps. This motivates our key question: Can improved credit assignment enhance RL training for LLMs? To address this, we propose VinePPO, a straightforward approach that leverages the flexibility of language environments to compute unbiased Monte Carlo-based estimates. Our method consistently outperforms PPO and other baselines across MATH and GSM8K datasets in less wall-clock time (up to 3.0x). Crucially, it achieves higher test accuracy for a given training accuracy, capturing more generalization signal per sample. These results emphasize the importance of accurate credit assignment in RL training of LLM.
Code available at https://github.com/McGill-NLP/VinePPO
RL for LLM, RLVR, Verifiable Rewards, Credit Assignment
@@ 1 Introduction
Reinforcement learning (RL) has become instrumental in training large language models (LLMs) to solve complex reasoning tasks such as mathematical problem solving (DeepSeek-AI et al., 2025), web navigation (Putta et al., 2024), or code generation (OpenAI, 2024). In these settings, LLMs often engage in extended reasoning steps, executing multiple actions to arrive at a solution. However, not all steps are equally impactful—some contribute significantly, while others are irrelevant or detrimental. For example, in Figure 1.a, only step $s_{2}$ provides a key insight. Indeed, most reasoning steps generated by a model do not affect the chance of it solving the problem (Figure 1.b). Identifying the contribution of each action is crucial for improving model performance. However, this is inherently difficult due to the significant delay between actions and their eventual effect. This issue, known as the Credit Assignment (CA) problem, is a core challenge in RL (Sutton & Barto, 1998).
@@
Proximal Policy Optimization (PPO; Schulman et al. 2017) addresses credit assignment through a value network (or critic), a mechanism retained in its application to RL-based finetuning of LLMs (Ouyang et al., 2022). This network, typically a separate model initialized from a pretrained checkpoint, is trained during PPO finetuning to estimate the expected cumulative rewards (or value) of an intermediate action. In Figure 1.b, an ideal value network would assign high value to step $s_{2}$ and subsequent steps, where the model has a high chance of successfully solving the problem. PPO uses these value estimates to measure the advantage of each action and update the model accordingly.
Prompt ($s_{0}$) $\hat{p}(\text{correct}|s_{:t})$    Advantage
@@
$0.4$ n/a Let $a$ and $b$ be nonzero real numbers such that $(2-7i)(a+bi)$ is pure imaginary. Find $\frac{a}{b}.$
Chain-of-Thought Response
$s_{1}$$s_{1}$$0.4$ $0.0$ We can expand the left-hand side to get
$s_{2}$$s_{2}$$\mathbf{1.0}$ $\mathbf{0.6}$ $(2-7i)(a+bi)=(2a+7b)+(-7a+2b)i.$
$s_{3}$$s_{3}$$1.0$ $0.0$ This is pure imaginary if and only if the real part is 00, i.e.
$s_{4}$$s_{4}$$1.0$ $0.0$ $02a+7b=$ 0 .
$s_{5}$$s_{5}$$1.0$ $0.0$ Then $a=-\frac{7}{2}b,$ so $\frac{a}{b}=-\frac{7}{2}.$
(a) A Sample Response
(b) Advantage
@@
Figure 1: (Left) A response generated by the model. The notation $\hat{p}(\text{correct}|s_{:t})$ represents the estimated probability of successfully solving the problem at step $t$. Here, only step $s_{2}$ is critical; after this, the model always completes the solution correctly. (Right) The distribution of advantages, defined as $\hat{p}(\text{correct}|s_{:t+1})-\hat{p}(\text{correct}|s_{:t})$, collected over a subset of MATH dataset (Hendrycks et al., 2021). Most steps show little or no advantage over the preceding step.
However, recent approaches such as DPO (Rafailov et al., 2023) or GRPO (Shao et al., 2024) simplify PPO’s design by discarding fine-grained credit assignment and treating all tokens equally. Despite such simplifications, they often demonstrate strong performance (Xu et al., 2024; Chang et al., 2023). This challenges classic RL principles, where accurate CA is considered critical for optimal performance (Sutton & Barto, 1998; Greensmith et al., 2001), especially in tasks with delayed rewards. In this work, we address this apparent discrepancy by showing that PPO’s credit assignment mechanism, the value network, performs poorly in practice. Our systematic evaluation (Section 7) on tasks requiring chain-of-thought reasoning reveals that the value network often provides imprecise estimates and fails to differentiate between promising and unproductive steps, which could explain why simplified approaches achieve comparable results without explicit fine-grained CA.
These findings motivate a central question: If we improve credit assignment in PPO rather than discarding it, how much can we enhance the RL training of LLMs? To explore this, we propose VinePPO (Figure 2), which computes unbiased value estimates of the intermediate states with Monte Carlo (MC) estimation instead of employing value networks. Our key insight is that language-based environments allow us to reset directly to any intermediate state simply by re-feeding the partial context, enabling efficient MC rollouts without the massive overhead usually seen in generic RL environments. VinePPO preserves PPO’s overall framework but addresses the CA challenge fundamentally.
We empirically evaluate the effectiveness and computational efficiency of MC value estimation in VinePPO. Across multiple mathematical reasoning tasks, VinePPO consistently outperforms PPO and other credit assignment-free baselines. While its per-iteration runtime is generally slower due to MC sampling, VinePPO surpasses the peak performance of baselines with fewer gradient steps and ultimately less wall-clock time. Importantly, VinePPO achieves higher test accuracy for a given training accuracy, capturing more generalization signal per fitted training sample. This is critical, as genuinely challenging verifiable reasoning tasks are scarce. These results underscore the importance of CA in RL-training of LLMs and highlight VinePPO as a straightforward alternative to value network-based approaches.
Our contributions are summarized as follows:
  • We analyze PPO’s value network in reasoning tasks and find it often misestimates intermediate values, barely outperforming a random chance in ranking candidate steps.
  • We propose VinePPO, leveraging the flexibility of language environments to compute unbiased, MC-based value estimates without relying on a separate critic.
@@
  • We empirically highlights the benefits of refined CA. VinePPO achieves the peak performance of baselines with less wall-clock time (up to 3.0x), better KL-divergence trade-off while exhibiting better generalization slope.
${\bm{x}}$$s_{1}$$s_{2}$$\tau_{2}$$\tau_{1}$$\tau_{3}$ RLOO: $\hat{V}_{\mathcal{R}}({\bm{x}})=\frac{1}{G-1}\sum_{2}^{G}R(\tau_{i})$ GRPO: $V_{\mathcal{G}}({\bm{x}})=\frac{1}{G}\sum_{1}^{G}R(\tau_{i})$ $b_{1}$$=\hat{V}_{\mathcal{R}}({\bm{x}})=\frac{1}{2}(0+1)$
  $=0.5$
$b_{2}$$=\hat{V}_{\mathcal{R}}({\bm{x}})=\frac{1}{2}(0+1)$
  $=0.5$
RLOO / GRPO
${\bm{x}}$$s_{1}$$s_{2}$ $\tau_{2}$$\tau_{1}$$\tau_{3}$ $b_{1}$$=\hat{V}_{\phi}(s_{1})$
  $=0.1934$
$b_{2}$$=\hat{V}_{\phi}(s_{2})$
  $=0.5733$
$\hat{V}_{\phi}(s_{t})=\mathrm{ValNet}(s_{t};\phi)$ PPO
${\bm{x}}$$s_{1}$$s_{2}$$\eta_{1}$$\eta_{1}$$\eta_{3}$$\eta_{1}^{\prime}$$\eta_{1}^{\prime}$$\eta_{3}^{\prime}$ $\tau_{2}$$\tau_{1}$$\tau_{3}$ $b_{1}$$=\hat{V}_{\mathrm{MC}}(s_{1})$
  $=0.66$
$b_{2}$$=\hat{V}_{\mathrm{MC}}(s_{2})$
  $=1.00$
$\hat{V}_{\mathrm{MC}}(s_{t})=\frac{1}{K}\sum_{k}R(\eta_{k})$ VinePPO
Figure 2: Comparison of credit assignment mechanisms applied on training trajectories $\tau_{i}$’s, depicted for states $s_{1}$ and $s_{2}$. (Left) RLOO and GRPO both treat all intermediate states equally and use the average return of trajectory group $\tau_{i}\sim\pi(\cdot|x)$ for the policy‐gradient baselines $b_{1}$ and $b_{2}$. GRPO additionally normalize these returns to have a unit variance. In the case of RLOO, the computed baseline could be viewed as MC estimate of value but solely for the initial state. (Middle) PPO trains a separate model to predict values for each state $s_{t}$. (Right) VinePPO generate auxiliary rollouts $\eta_{k}\sim\pi(\cdot|s_{t})$ to obtain MC estimate of state $s_{t}$’s value. Note that $\eta_{k}$’s are only used for value estimation—not to update the policy directly.
@@ 2 Related Work
PPO, as applied in RL from Human Feedback (RLHF, Ouyang et al. 2022), pioneered RL finetuning of LLMs. However, its computational overhead and hyperparameter sensitivity led to the development of simpler alternatives. RL-free methods such as DPO (Rafailov et al., 2023) operate in a bandit setting, treating the entire response as a single action. Similarly, rejection sampling methods like RestEM (Singh et al., 2024) finetune on full high-reward responses. RLOO (Ahmadian et al., 2024) and GRPO (Shao et al., 2024) abandon PPO’s value network, instead using average reward from multiple samples as a policy gradient baseline. Recent work has emphasized finer credit assignment, with Hwang et al. (2024) and Setlur et al. (2024) introducing MC-based methods to detect key errors in reasoning chains for use as ad-hoc mechanisms in DPO. Our work, by contrast, fully embraces the RL training, with the target of fixing CA in principle. Parallel efforts have also focused on building better verifiers and reward models for per-step feedback, with recent attempts to automate their data collection using MC rollouts (Ma et al., 2023; Uesato et al., 2022; Luo et al., 2024; Wang et al., 2024). Our method is orthogonal to these methods, operating within PPO-based training to optimize a given reward, instead of designing new ones.
@@
Deep RL algorithms are typically categorized into value-based and policy-based methods. Policy-based methods like PPO usually employ critic networks for value prediction. An exception is the “Vine” variant of TRPO (Schulman et al., 2015), which uses MC samples for state value estimation. The authors, however, note that the Vine variant is limited to environments that allow intermediate state resets, rare in typical RL settings [^1] This is reflected in the design of Gym (Towers et al., 2024), which only allows resets to the initial state.. However, language generation – when formulated as RL environment – enables such intermediate reset capabilities. In domains with similar reset capabilities, such as Go and Chess, MC-based methods like AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) have emerged. AlphaGo’s architecture includes a policy, trained using expert moves and self-play, and a value network that predicts game outcomes. At inference, it employs tree search guided by MC rollouts and value network to select optimal moves. AlphaZero advances this approach by distilling MCTS outcomes into the policy. Recent works have adapted AlphaZero’s principles to LLMs, employing similar search techniques for inference and trajectory distillation (Xie et al., 2024; Chen et al., 2024; Wan et al., 2024; Zhang et al., 2024; Hao et al., 2023). While this is a promising direction, our method is not an MCTS approach; it uses MC samples solely for value estimation during PPO training to improve credit assignment.
Figure 3: Distribution of predicted values for each state vs. ground truth (see Section 7 for details) for DeepSeekMath 7B on the MATH dataset, highlighting the nature of errors: PPO exhibits biased value predictions, whereas VinePPO remains unbiased. Note that RLOO/GRPO do not predict values; we plot their computed baselines against the ground truth value solely for demonstration.
@@ 3 Background
We focus on the RL tuning phase, following Ouyang et al. (2022); Shao et al. (2024). In this section, we provide an overview of actor-critic finetuning as implemented in the standard PPO framework.
@@
where $\mathcal{D}$ is the dataset of prompts, $\mathcal{R}({\bm{x}};{\bm{y}})$ is the sequence-level reward function, and $\beta$ controls the strength of the KL penalty. Note that the policy $\pi_{\theta}$ is initialized from $\pi_{\mathrm{ref}}$.
@@
${\bm{g}}_{\mathrm{pg}}=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=0}^{T-1}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})A(s_{t},a_{t})\right],$ (2)
where $s_{t}={\bm{x}};{\bm{y}}_{<t},\ a_{t}=y_{t}$, and $A(s_{t},a_{t})$ is the advantage function. If $A(s_{t},a_{t})>0$, ${\bm{g}}_{\mathrm{pg}}$ will increase the probability of action $a_{t}$ in state $s_{t}$, and decrease it when $A(s_{t},a_{t})<0$. Intuitively, the advantage function quantifies how much better action $a_{t}$ is compared to average actions taken in state $s_{t}$ under the policy. Formally, it is defined as:
$A(s_{t},a_{t})$ $=Q(s_{t},a_{t})-V(s_{t})$
$=r_{t}+\gamma V(s_{t+1})-V(s_{t}).$ (3)
@@
In practice, the advantage $A(s_{t},a_{t})$ is not known beforehand and is typically estimated by first using a value network $\hat{V}_{\phi}$ to approximate the true value function $V(s_{t})$, then substituting the learned values into Section 3 or alternative methods like GAE (Schulman et al., 2016). The value network is parameterized by $\phi$ and trained alongside the policy network $\pi_{\theta}$. The training objective for the value network minimizes the mean squared error between the predicted value and the empirical return:
@@
where $G_{t}=\sum_{t^{\prime}=t}^{T-1}r_{t^{\prime}}$ is the empirical return from state $s_{t}$. PPO uses the same objective for $\hat{V}_{\phi}$ but applies clipping for training stability (see Section A.1 for details). In RL-tuning of LLMs, the value network is often initialized using the initial SFT policy $\pi_{\mathrm{ref}}$ (or the reward model when available), with the language modeling head swapped out for a scalar head to predict values (Zheng et al., 2023). This setup leverages the prior knowledge of the pretrained model.